OpenStudy (anonymous):

Assume Kellogg's produced 715,000 boxes of corn flakes this year. This was 110% of the annual production last year. What was last year's annual production?

OpenStudy (anonymous):

y = 715,000/ 1.1 = 650,000 boxes

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes a little but can you show me how you work that out

OpenStudy (anonymous):

110% of last year = 715,000

OpenStudy (anonymous):

-> Last year = 715,000/ 110%

OpenStudy (anonymous):

ok thanks